# Beyond Linear Approximations: A Novel Pruning Approach for Attention Matrix

### v1

@@ 7 Conclusion
This paper introduced a novel approach to LLM weight pruning that directly optimizes for approximating the attention matrix. We provided theoretical guarantees for the convergence of our Gradient Descent-based algorithm to a near-optimal pruning mask solution. Preliminary results demonstrated the method’s effectiveness in maintaining model performance while reducing computational costs. This work establishes a new theoretical foundation for pruning algorithm design in LLMs, potentially enabling more efficient inference on resource-constrained devices.
@@ Acknowledgement
The appendix is organized as follows. In Section A, we give the preliminary of our paper. In Section B, we provide detailed gradient analysis of loss function. In Section C, we provide details about how we integrate the gradient of loss function into matrix form. In Section D, we bound some basic functions to be used later. In Section E, we provide proof for the Lipschitz property of the gradient of the loss function. In Section F, we provide proof of convergence for GD.
Appendix A Preliminary
In Section A.1, we introduce some notations we use in this paper. In Section A.2, we provide some basic facts.
A.1 Notations
@@
A.2 Facts
Fact A.1 (Indexing).
@@ A.2 Facts
Fact A.4 (Basic algebra).
@@
Fact A.5 (Norm bounds).
@@
where the first step, the second step and the third step follow from Fact A.3, the fourth step follows from Fact A.4.
Appendix B Gradient Calculation
B.1 Definitions
@@
Definition B.1 (Exponential function $u$, $\widetilde{u}$).
If the following condition hold
@@
Definition B.2 (Sum function of softmax $\alpha$, $\widetilde{\alpha}$).
If the following condition hold
@@
  • Let $u$, $\widetilde{u}(M)$ be defined as Definition B.1.
@@
Definition B.3 (Softmax probability function $f$, $\widetilde{f}$).
@@
  • Let $u$, $\widetilde{u}(M)$ be defined as Definition B.1.
  • Let $\alpha$, $\widetilde{\alpha}(M)$ be defined as Definition B.2.
@@
Then, we introduce the one unit loss function.
Definition B.4 (One unit loss function $c$).
@@
  • Let $f$, $\widetilde{f}$ be defined in Definition B.3.
@@
Definition B.5 (Reconstruction Error ${\cal L}_{\mathrm{attn}}$).
@@
Definition B.6 (Regularization Term ${\cal L}_{\mathrm{reg}}$).
@@
Definition B.7 (Overall loss function ${\cal L}$).
@@
  • Let ${\cal L}_{\mathrm{attn}}(M)$ be defined in Definition B.5.
  • Let ${\cal L}_{\mathrm{reg}}(M)$ be defined in Definition B.6.
@@
B.2 Gradient for Each Row of $X(M\circ W)X^{\top}$
@@
where the first and second step follows from Fact A.1, the third step follows from Fact A.3.
We further compute Eq. (B.2):
@@
where the first follows from Fact A.4, the second step follows from for any matrix $X$, $X_{i,j}=(X^{\top})_{j,i}$, the third step follows from Fact A.7, and the fourth step follows from for any matrix $X$, $X_{i,j}=(X^{\top})_{j,i}$.
@@
where the first step follows from Eq. (B.2) and Eq. (B.2), and the second step and the third step follows from basic algebra. ∎
We introduce the Lemma of gradient for each row of $\widetilde{u}(M)$.
B.3 Gradient for Each Row of $\widetilde{u}(M)$
Lemma B.9.
If the following condition hold:
  • Let $\widetilde{u}(M)$ be defined in Definition B.1.
@@ B.3 Gradient for Each Row of $\widetilde{u}(M)$
where the first step follows from Definition B.1, the second step follows from Fact A.3, the third step follows from Definition B.1, and the fourth step follows from Lemma B.8. ∎
B.4 Gradient for Each Entry of $\widetilde{\alpha}(M)$
@@
  • Let $\widetilde{u}(M)$ be defined in Definition B.1.
  • Let $\widetilde{\alpha}(M)$ be defined in Definition B.2.
@@
where the first step follows from Definition B.2, the second step follows from product rule of inner product in Fact A.3, the third step follows from product rule of Hadamard product in Fact A.3, the fourth step follows from Lemma B.9, and the last step follows from Fact A.4. ∎
B.5 Gradient for Each Entry of $\widetilde{f}(M)$
We introduce the Lemma of gradient for each entry of $\widetilde{f}(M)$.
Lemma B.11.
@@ B.5 Gradient for Each Entry of $\widetilde{f}(M)$
  • Let $\widetilde{u}(M)$ be defined in Definition B.1.
  • Let $\widetilde{\alpha}(M)$ be defined in Definition B.2.
  • Let $\widetilde{f}(M)$ be defined in Definition B.3.
@@
where the first step follows from Definition B.3, and the second step follows from Fact A.3.
@@
where the first step follows from Fact A.3, the second step follows from Lemma B.10, the third step follows from basic algebra, and the fourth step follows from Definition B.3.
@@
where the first step and the second step follow from basic algebra, the third step follows from Lemma B.9, and the fourth step follows from Definition B.3.
@@
where the first step follows from Eq. (B.5), and the second step follows from Eq. (B.5) and Eq. (B.5). ∎
B.6 Gradient for Each Entry of c(M)
@@
  • Let $\widetilde{f}(M)$, $f$ be defined in Definition B.3.
  • Let $c(M)$ be defined in Definition B.4.
@@
where the first step follows from Definition B.4, the second step follows from Fact A.3, and the third step follows from Lemma B.11. ∎
B.7 Gradient for ${\cal L}_{\mathrm{attn}}(M)$
We introduce the Lemma of gradient for ${\cal L}_{\mathrm{attn}}(M)$.
Lemma B.13.
@@ B.7 Gradient for ${\cal L}_{\mathrm{attn}}(M)$
  • Let $\widetilde{f}(M)$ be defined in Definition B.3.
  • Let $c(M)$ be defined in Definition B.4.
  • Let ${\cal L}_{\mathrm{attn}}(M)$ be defined in Definition B.5.
@@
where the first step follows from Definition B.5, the second step follows from the definition of Frobenius norm of matrix, the third step follows from Fact A.3, and the fourth step follows from Fact A.3.
Following Lemma B.12, we have
@@
B.8 Gradient for ${\cal L}_{\mathrm{reg}}(M)$
We introduce the Lemma of gradient for ${\cal L}_{\mathrm{reg}}(M)$.
Lemma B.14.
If the following condition hold:
  • Let ${\cal L}_{\mathrm{reg}}(M)$ be defined in Definition B.6.
@@ B.8 Gradient for ${\cal L}_{\mathrm{reg}}(M)$
where we have definition:
@@
where the first step follows from Definition B.6, the second step follows from the definition of Frobenius norm of matrix, and the third step follows from Fact A.3. ∎
B.9 Gradient for ${\cal L}(M)$
We introduce the Lemma of gradient for ${\cal L}(M)$.
Lemma B.15.
@@ B.9 Gradient for ${\cal L}(M)$
  • Let $\widetilde{u}(M)$ be defined in Definition B.1.
  • Let $\widetilde{\alpha}(M)$ be defined in Definition B.2.
  • Let $\widetilde{f}(M)$ be defined in Definition B.3.
  • Let ${\cal L}_{\mathrm{attn}}(M)$ be defined in Definition B.5.
  • Let ${\cal L}_{\mathrm{reg}}(M)$ be defined in Definition B.6.
  • Let ${\cal L}(M)$ be defined in Definition B.7.
@@
where the first step follows from Definition B.7, the second step follows from Fact A.3, and the third step follows from Lemma B.13 and Lemma B.14. ∎
Appendix C Matrix Form
C.1 Matrix Form of $B(M)$
@@
  • Let $c(M)$ be defined in Definition B.4.
  • Let $\widetilde{f}(M)$ be defined in Definition B.3.
@@
Lemma C.2 (Matrix view of $B_{1}(M)$).
@@
  • Let $B_{1}(M,i_{1},j_{1}):=c(M)_{i_{0},j_{0}}\widetilde{f}(M)_{i_{0},j_{0}}W_{i_{1},j_{1}}X_{i_{0},i_{1}}X_{j_{0},j_{1}}$, which is defined in Lemma B.15
@@
where the first step follows from the definition of $C_{1}$, the second step follows from Fact A.1, and the third step follows from Fact A.2.
Following from Fact A.1, we can get $j_{1}$-th column of $C_{1}$
@@
where the second step follows from Fact A.2.
Following from Fact A.1, we can get $C_{1}(M)$
@@
where the second step follows from Fact A.4.
@@
where the first step follows from Eq. (C.1), the second step follows from basic algebra, the third step follows from Fact A.4, and the fourth step follows from Definition C.1.
Then following from Fact A.2, we have
@@
Lemma C.3 (Matrix view of $B_{2}(M)$).
@@
  • Let $B_{2}(M,i_{1},j_{1}):=-c(M)_{i_{0},j_{0}}\widetilde{f}(M)_{i_{0},j_{0}}\langle\widetilde{f}(M)_{i_{0}},W_{i_{1},j_{1}}X_{i_{0},i_{1}}X_{*,j_{1}}\rangle$ be defined in Lemma B.15.
@@
where the first step follows from the definition of $C_{2}$, the second step, the third step and the fourth step follows from Fact A.4.
Following from Fact A.1, we can get $j_{1}$-th column of $C_{2}$
@@
where the second step and the fourth step follows from Fact A.4, and the third step follows from Fact A.1.
Following from Fact A.1, we can get $C_{2}$.
@@
where the first step follows from Eq. (7), the second step and the third step follows from Fact A.4.
Following from Fact A.2, we have
@@
where the third step follows from Definition C.1. ∎
@@
Lemma C.4 (Matrix view of $B_{3}(M)$).
@@
  • Let $B_{3}(M,i_{1},j_{1}):=\lambda M_{i_{1},j_{1}}$ be defined in Lemma B.15.
@@
C.2 Matrix Form of $\frac{\mathrm{d}}{\mathrm{d}M}{\cal L}(M)$
We introduce the matrix form of overall loss function.
Theorem C.5 (Close form of gradient, formal version of Theorem 5.3).

### v2

@@ 7 Conclusion
This paper introduces a novel approach to LLM weight pruning that directly optimizes for approximating the attention matrix. We provide theoretical guarantees for the convergence of our Gradient Descent-based algorithm to a near-optimal pruning mask solution. Experimental results demonstrated the method’s effectiveness in maintaining model performance while reducing computational costs. This work establishes a new theoretical foundation for pruning algorithm design in LLMs, potentially enabling more efficient inference on resource-constrained devices.
@@ Acknowledgement
The appendix is organized as follows. In Section A, we reviewed more literature related to our paper. In Section B, we give the preliminary of our paper. In Section C, we provide a detailed gradient analysis of the loss function. In Section D, we provide details about how we integrate the gradient of loss function into matrix form. In Section E, we bound some basic functions to be used later. In Section F, we provide proof for the Lipschitz property of the gradient of the loss function. In Section G, we provide proof of convergence for GD.
Appendix A More Related Work
Large Language Models.
Transformer-based neural networks [136] have rapidly emerged as the dominant architecture for natural language processing in machine learning. When expanded to billions of parameters and trained on vast, diverse datasets, these systems are typically termed large language models (LLMs) or foundation models [14]. Prominent LLM examples encompass BERT [30], PaLM [26], Llama [133, 102], and GPT4o [109], which display adaptable competencies [13] across numerous downstream applications. To enhance LLMs for domain-specific uses, researchers have created multiple adaptation approaches. These include: adapter modules [59, 47, 159, 118]; calibration mechanisms [163, 166]; multitask refinement [46, 135, 150, 149]; along with prompt engineering [77], scratchpad approaches [108], instruction optimization [16, 80, 103], symbolic adaptation [138, 151, 152], black-box adjustments [122], human-aligned reinforcement learning [111], and structured reasoning techniques [76, 154, 142, 165]. Contemporary investigations cover tensor architecture innovations [9, 97, 119, 164], efficiency enhancements [7, 18, 25, 53, 54, 56, 63, 64, 73, 85, 83, 84, 90, 88, 95, 121, 127, 128, 132, 137, 139, 145], plus ancillary studies [21, 33, 17, 35, 48, 50, 86, 87, 96, 99, 125, 126, 131, 134, 147, 148, 157, 167, 24, 81, 72, 93, 22, 23, 49, 62, 60, 141].
Appendix B Preliminary
In Section B.1, we introduce some notations we use in this paper. In Section B.2, we provide some basic facts.
B.1 Notations
@@
B.2 Facts
Fact B.1 (Indexing).
@@ B.2 Facts
Fact B.4 (Basic algebra).
@@
Fact B.5 (Norm bounds).
@@
where the first step, the second step, and the third step follow from Fact B.3, the fourth step follows from Fact B.4.
Appendix C Gradient Calculation
C.1 Definitions
@@
Definition C.1 (Exponential function $u$, $\widetilde{u}$).
If the following conditions hold
@@
Definition C.2 (Sum function of softmax $\alpha$, $\widetilde{\alpha}$).
If the following conditions hold
@@
  • Let $u$, $\widetilde{u}(M)$ be defined as Definition C.1.
@@
Definition C.3 (Softmax probability function $f$, $\widetilde{f}$).
@@
  • Let $u$, $\widetilde{u}(M)$ be defined as Definition C.1.
  • Let $\alpha$, $\widetilde{\alpha}(M)$ be defined as Definition C.2.
@@
Then, we introduce the one-unit loss function.
Definition C.4 (One unit loss function $c$).
@@
  • Let $f$, $\widetilde{f}$ be defined in Definition C.3.
@@
Definition C.5 (Reconstruction Error ${\cal L}_{\mathrm{attn}}$).
@@
Definition C.6 (Regularization Term ${\cal L}_{\mathrm{reg}}$).
@@
Definition C.7 (Overall loss function ${\cal L}$).
@@
  • Let ${\cal L}_{\mathrm{attn}}(M)$ be defined in Definition C.5.
  • Let ${\cal L}_{\mathrm{reg}}(M)$ be defined in Definition C.6.
@@
C.2 Gradient for Each Row of $X(M\circ W)X^{\top}$
@@
where the first and second step follows from Fact B.1, the third step follows from Fact B.3.
We further compute Eq. (C.2):
@@
where the first follows from Fact B.4, the second step follows from for any matrix $X$, $X_{i,j}=(X^{\top})_{j,i}$, the third step follows from Fact B.7, and the fourth step follows from for any matrix $X$, $X_{i,j}=(X^{\top})_{j,i}$.
@@
where the first step follows from Eq. (C.2) and Eq. (C.2), and the second step and the third step follow from basic algebra. ∎
We introduce the Lemma of the gradient for each row of $\widetilde{u}(M)$.
C.3 Gradient for Each Row of $\widetilde{u}(M)$
Lemma C.9.
If the following conditions hold:
  • Let $\widetilde{u}(M)$ be defined in Definition C.1.
@@ C.3 Gradient for Each Row of $\widetilde{u}(M)$
where the first step follows from Definition C.1, the second step follows from Fact B.3, the third step follows from Definition C.1, and the fourth step follows from Lemma C.8. ∎
C.4 Gradient for Each Entry of $\widetilde{\alpha}(M)$
@@
  • Let $\widetilde{u}(M)$ be defined in Definition C.1.
  • Let $\widetilde{\alpha}(M)$ be defined in Definition C.2.
@@
where the first step follows from Definition C.2, the second step follows from the product rule of inner product in Fact B.3, the third step follows from the product rule of Hadamard product in Fact B.3, the fourth step follows from Lemma C.9, and the last step follows from Fact B.4. ∎
C.5 Gradient for Each Entry of $\widetilde{f}(M)$
We introduce the Lemma of the gradient for each entry of $\widetilde{f}(M)$.
Lemma C.11.
@@ C.5 Gradient for Each Entry of $\widetilde{f}(M)$
  • Let $\widetilde{u}(M)$ be defined in Definition C.1.
  • Let $\widetilde{\alpha}(M)$ be defined in Definition C.2.
  • Let $\widetilde{f}(M)$ be defined in Definition C.3.
@@
where the first step follows from Definition C.3, and the second step follows from Fact B.3.
@@
where the first step follows from Fact B.3, the second step follows from Lemma C.10, the third step follows from basic algebra, and the fourth step follows from Definition C.3.
@@
where the first step and the second step follow from basic algebra, the third step follows from Lemma C.9, and the fourth step follows from Definition C.3.
@@
where the first step follows from Eq. (C.5), and the second step follows from Eq. (C.5) and Eq. (C.5). ∎
C.6 Gradient for Each Entry of c(M)
@@
  • Let $\widetilde{f}(M)$, $f$ be defined in Definition C.3.
  • Let $c(M)$ be defined in Definition C.4.
@@
where the first step follows from Definition C.4, the second step follows from Fact B.3, and the third step follows from Lemma C.11. ∎
C.7 Gradient for ${\cal L}_{\mathrm{attn}}(M)$
We introduce the Lemma of the gradient for ${\cal L}_{\mathrm{attn}}(M)$.
Lemma C.13.
@@ C.7 Gradient for ${\cal L}_{\mathrm{attn}}(M)$
  • Let $\widetilde{f}(M)$ be defined in Definition C.3.
  • Let $c(M)$ be defined in Definition C.4.
  • Let ${\cal L}_{\mathrm{attn}}(M)$ be defined in Definition C.5.
@@
where the first step follows from Definition C.5, the second step follows from the definition of Frobenius’s norm of the matrix, the third step follows from Fact B.3, and the fourth step follows from Fact B.3.
Following Lemma C.12, we have
@@
C.8 Gradient for ${\cal L}_{\mathrm{reg}}(M)$
We introduce the Lemma of the gradient for ${\cal L}_{\mathrm{reg}}(M)$.
Lemma C.14.
If the following conditions hold:
  • Let ${\cal L}_{\mathrm{reg}}(M)$ be defined in Definition C.6.
@@ C.8 Gradient for ${\cal L}_{\mathrm{reg}}(M)$
where we have the definition:
@@
where the first step follows from Definition C.6, the second step follows from the definition of Frobenius’s norm of the matrix, and the third step follows from Fact B.3. ∎
C.9 Gradient for ${\cal L}(M)$
We introduce the Lemma of the gradient for ${\cal L}(M)$.
Lemma C.15.
@@ C.9 Gradient for ${\cal L}(M)$
  • Let $\widetilde{u}(M)$ be defined in Definition C.1.
  • Let $\widetilde{\alpha}(M)$ be defined in Definition C.2.
  • Let $\widetilde{f}(M)$ be defined in Definition C.3.
  • Let ${\cal L}_{\mathrm{attn}}(M)$ be defined in Definition C.5.
  • Let ${\cal L}_{\mathrm{reg}}(M)$ be defined in Definition C.6.
  • Let ${\cal L}(M)$ be defined in Definition C.7.
@@
where the first step follows from Definition C.7, the second step follows from Fact B.3, and the third step follows from Lemma C.13 and Lemma C.14. ∎
Appendix D Matrix Form
D.1 Matrix Form of $B(M)$
@@
  • Let $c(M)$ be defined in Definition C.4.
  • Let $\widetilde{f}(M)$ be defined in Definition C.3.
@@
Lemma D.2 (Matrix view of $B_{1}(M)$).
@@
  • Let $B_{1}(M,i_{1},j_{1}):=c(M)_{i_{0},j_{0}}\widetilde{f}(M)_{i_{0},j_{0}}W_{i_{1},j_{1}}X_{i_{0},i_{1}}X_{j_{0},j_{1}}$, which is defined in Lemma C.15
@@
where the first step follows from the definition of $C_{1}$, the second step follows from Fact B.1, and the third step follows from Fact B.2.
Following from Fact B.1, we can get $j_{1}$-th column of $C_{1}$
@@
where the second step follows from Fact B.2.
Following from Fact B.1, we can get $C_{1}(M)$
@@
where the second step follows from Fact B.4.
@@
where the first step follows from Eq. (D.1), the second step follows from basic algebra, the third step follows from Fact B.4, and the fourth step follows from Definition D.1.
Then following from Fact B.2, we have
@@
Lemma D.3 (Matrix view of $B_{2}(M)$).
@@
  • Let $B_{2}(M,i_{1},j_{1}):=-c(M)_{i_{0},j_{0}}\widetilde{f}(M)_{i_{0},j_{0}}\langle\widetilde{f}(M)_{i_{0}},W_{i_{1},j_{1}}X_{i_{0},i_{1}}X_{*,j_{1}}\rangle$ be defined in Lemma C.15.
@@
where the first step follows from the definition of $C_{2}$, the second step, the third step, and the fourth step follow from Fact B.4.
Following from Fact B.1, we can get $j_{1}$-th column of $C_{2}$
@@
where the second step and the fourth step follows from Fact B.4, and the third step follows from Fact B.1.
Following from Fact B.1, we can get $C_{2}$.
@@
where the first step follows from Eq. (8), the second step and the third step follow from Fact B.4.
Following from Fact B.2, we have
@@
where the third step follows from Definition D.1. ∎
@@
Lemma D.4 (Matrix view of $B_{3}(M)$).
@@
  • Let $B_{3}(M,i_{1},j_{1}):=\lambda M_{i_{1},j_{1}}$ be defined in Lemma C.15.
@@
D.2 Matrix Form of $\frac{\mathrm{d}}{\mathrm{d}M}{\cal L}(M)$
We introduce the matrix form of the overall loss function.
Theorem D.5 (Close form of gradient, formal version of Theorem 5.3).